Problem:
The length of a rectangular poool is 40 feet. the width is 24 feet. A walkway of uniform width surrounds the pool. The combined area of the pool and walkway is 2240 sqft. Find the width of the walkway?
Solution:

40s+40s+24s+24s = 2240
128s = 2240
s = 17.5
I dont know how to go about it... I know this is wrong!!!!
Plz help!!!! I do not have the ISBN because that is from one of my exams .

Answer by checkley75(3666)

40*24=960 SQUARE FEET FOR THE POOL.
(40+2X)(24+2X)=2240 THESE ARE THE NEW DIMENTIONS WHEN YOU ADD THE WALKWAY WITH A WIDTH OF X. tHE BEST WAY TO VISUALIZE THIS IS TO DRAW A DIAGRAM OF THE POOL AND ADD THE WALKWAY OF WIDTH X.
960+48X+80X+4X^2=2240
4X^2+128X+960-2240=0
4X^2+128X-1280=0
4(X^2+32X-320)=0
4(X-8)(X+40)=0
X-8=0
X=8 FEET ANSWER FOR THE WIDTH OF THE WALKWAY.
PROOF
(40+2*8)(24+2*8)=2240
(40+16)(24+16)=2240
56*40=2240
2240=2240